# On the Sylvester program and Cayley algorithm for vector partition reduction

### v1

@@ Abstract
A vector (compound) partition problem asks for a number of nonnegative integer solutions to a system of several linear Diophantine equations with integer nonnegative coefficients. J.J. Sylvester put forward an idea of reduction of vector partition to a sum of regular (scalar) partitions. For the simplest case of two equations with positive coefficients A. Cayley performed a reduction of the corresponding double partition to a sum of scalar partitions using an algorithm subject to a set of conditions on the coefficients. We suggested [7] a modification of the original Cayley algorithm for the cases when these conditions are not satisfied. This result is extended to arbitrary number of the Diophantine equations to accomplish the Sylvester program of the vector partition reduction to a combination of scalar partitions.
@@ 1 Introduction
The summation is made over all prime roots of unity $\rho_{j}=\exp(2\pi in/j)$ for $n$ relatively prime to $j$ (including unity) and smaller than $j$. It is possible to express any Sylvester wave $W_{j}(s,{\bf d})$ and SPF $W(s,{\bf d})$ as a finite sum of the Bernoulli polynomials of higher order [4].
@@
Several approaches were suggested for VPF computation including the method of residues [1, 11] and geometric decomposition into so called chambers [8] – regions in $l$-dimensional space where the partition function is characterized by a specific expression. A method of VPF computation was suggested being a direct generalization of the approach developed in [4] where vector Bernoulli and Eulerian polynomials of higher order were introduced to find explicit expression for $W({\bf s},{\bf D})$ [5]. A drawback of this approach is that it does not provide any mechanism to define chamber boundaries. As SPF computation requires just well-known functions [4] it is promising to obtain a reduction method expressing a vector partition through a combination of scalar ones.
The first step in this direction was made in the 19th century by J.J. Sylvester [9, 10] who suggested an iterative procedure of reduction of a vector partition into a sum of SPFs; in [9] he gave an example of such a reduction for a double partition. Based on this approach A. Cayley developed [2] an algorithm of double partition reduction and established a set of conditions of this method applicability. Cayley showed that each column ${\bf c}_{i}$ of the two-row matrix ${\bf D}$ gives rise to a scalar partition and elements of this column must be relatively prime (we call it P-column). Another limitation of the Cayley method is that the columns ${\bf c}_{i}$ should represent noncollinear vectors. A short historical excurse into the (seemingly forgotten) ideas of Sylvester and Cayley is given in Section 2 while the essence of the original Cayley algorithm is presented in Section 3.1.
@@
In Section 4 we extend the reduction algorithm to $l$-tuple (for arbitrary $l\geq 2$) vector partition defined by a $l\times m$ nonnegative generator matrix ${\bf D}$ made of $m$ columns ${\bf c}_{i}$. The first step of the algorithm discussed in Section 4.1 is the partial fraction expansion of the generating function (6) into $m$ terms corresponding to each column. The contribution of a column is computed in Section 4.2 and it is either a single partition for P-columns (Section 4.2.1) or a superposition of such partitions (Section 4.2.2) for NP-column. Two examples of triple partition reduction are given in Appendices.
The case of collinear columns of the generator matrix ${\bf D}$ is addressed in Section 5 and we show that in this case a vector partition reduces to a convolution of scalar partitions. We also demonstrate that the reduction procedure does not generate new collinear columns.
Other details of computational implementation of the reduction step are presented in Sections 6.1 and 6.2 for P-column and NP-column respectively. We show in Section 6.3 that the reduction of a triple partition always produces NP-columns. This implies that the reduction of any VPF for a matrix with $l>2$ rows inevitably leads to application of the extended version of the Cayley algorithm. Additional details of the iterative reduction procedure are discussed in Section 6.4 while the concluding remarks are presented in Section 7.
@@ 2 Sylvester-Cayley method
Both manuscripts [9] and [2] were mentioned by L.E. Dickson in the then comprehensive volume [3] on the theory of numbers devoted to Diophantine analysis, but the literature search shows that both the reduction idea of Sylvester and the result of Cayley were forgotten. There are no publications discussing Sylvester research program of VPF reduction to SPFs. In my opinion the ideas and results of Sylvester and Cayley deserve not to be neglected, and this manuscript is an attempt to revive them and pave way to a general algorithm for vector partition expression through scalar partitions only.
@@ 3.4 Zero in last row of matrix ${\bf D}$
The matrix $\bm{D}^{\prime}_{1}$ used in derivation of (30) is assumed to be free of zeros in its last row. When this matrix has another column ${\bf c}_{j}$ with second zero element the columns ${\bf c}_{1}$ and ${\bf c}_{j}$ are collinear and we refer the reader to Section 5 that considers the general case of collinear columns.
@@ 4.2.1 Compact form
It should be reminded here that the final goal of the Sylvester program is the VPF reduction to a set of SPFs and thus the elimination of a single (last) row of $l\times m$ matrix is just the first step of the program. We will discuss the computation aspects of the Sylvester program further in Section 6.4 but now we have to underline that the reduction procedure described above is valid only when the matrix under reduction has last row void of nonpositive elements. It follows from (50) that some columns ${\bf c}^{1}_{j}$ might have negative elements $c^{1}_{jl}<0$ in the last position.
@@ 5 Vector partition with collinear columns
5.1 Reduction to SPF convolution
Call an integer generator matrix competent if its columns (viewed as vectors) are noncollinear and incompetent otherwise. Assume that vectors corresponding to the first $n$ columns of the incompetent generator matrix D are parallel to each other and rewrite the linear system ${\bf D}\cdot{\bf x}={\bf s}$ as follows
@@ 5.1 Reduction to SPF convolution
where $u_{j}$ are positive integers. The problem (66) is equivalent to
${\bf s}-k{\bf c}=\sum_{i=n+1}^{m}{\bf c}_{i}x_{i},\quad k=\sum_{i=1}^{n}u_{i}x_{i},\quad 0\leq k\leq k_{max}=\min(s_{r}/c_{r}),\quad 1\leq r\leq l.$ (67)
@@
Introduce a vector ${\bf u}=\{u_{1},u_{2},\ldots,u_{n}\}$ and a competent matrix with $(m-n)$ columns ${\bf D}_{n+1}=\{{\bf c}_{n+1},\ldots,{\bf c}_{m}\},$ for which the corresponding double partition $W({\bf D}_{n+1},{\bf s}-k{\bf c})$ admits a reduction to SPFs either as a sum of $W_{i}^{l}$ or a mixture of $W^{l}_{i}$ and $\bar{W}^{l}_{i}$. The number of solutions of the system (67) for given value of $k$ is equal to a number of solutions of the first equation given by the partition $W({\bf D}_{n+1},{\bf s}-k{\bf c})$ multiplied by a number of solutions of the second equation, i.e., by a scalar partition $W(k,{\bf u})$. Then the vector partition $W({\bf D},{\bf s})$ is equivalent to a convolution
$W({\bf D},{\bf s})=\sum_{k=0}^{k_{max}}W(k,{\bf u})W({\bf D}_{n+1},{\bf s}-k{\bf c}),\quad k_{max}=\min(s_{i}/c_{i}),\quad 1\leq i\leq l.$ (68)
of the scalar partition and the vector one built on the competent matrix ${\bf D}_{n+1}$. This partition itself allows reduction using the algorithm described in Section 4. As the final goal is a reduction of a vector partition through a combination of SPFs the reduction algorithm should be applied $(l-1)$ times to reach that goal. It is instructive to know whether the repeated application of the reduction step to partition based on a competent matrix might produce incompetent matrices.
5.2 Reduction process preserves column noncollinearity
[…]
${\cal D}^{lk}_{1r}=c_{1l}c_{rk}-c_{1k}c_{rl},\quad{\cal D}^{lk}_{1p}=c_{1l}c_{pk}-c_{1k}c_{pl}.$
If these two columns $\tilde{\bf c}_{r},\tilde{\bf c}_{p}$ are collinear their wedge product $\bm{W}=\tilde{\bf c}_{r}\wedge\tilde{\bf c}_{p}=0$ vanishes. The wedge product $\bm{W}$ is a $(l-1)\times(l-1)$ matrix and its element $W_{km}$ is given by
$W_{km}={\cal D}^{lk}_{1r}{\cal D}^{lm}_{1p}-{\cal D}^{lk}_{1p}{\cal D}^{lm}_{1r}.$
Direct computation shows that
$W_{km}=c_{1l}{\cal D}^{lkm}_{1rp},\quad\quad{\cal D}^{lkm}_{1rp}=\begin{vmatrix}c_{1k}&c_{rk}&c_{pk}\\ c_{1m}&c_{rm}&c_{pm}\\ c_{1l}&c_{rl}&c_{pl}\end{vmatrix},$
is a product of $3\times 3$ determinant ${\cal D}^{lkm}_{1rp}$ built on the columns $\{{\bf c}_{1},{\bf c}_{r},{\bf c}_{p}\}$ of the initial matrix ${\bf E}$ and the element $c_{1l}$. Collinearity of $\tilde{\bf c}_{r}$ and $\tilde{\bf c}_{p}$ requires that all $W_{km}=0$ and it is possible when either $c_{1l}=0$ or all determinants ${\cal D}^{lkm}_{1rp}$ vanish simultaneously. The condition $c_{1l}=0$ represents the case of a column with last zero element and this column does not contribute to partition reduction. The competence of the generator matrix ${\bf D}$ prevents satisfaction of ${\cal D}^{lkm}_{1rp}=0$ and thus the $(l-1)$-row matrix ${\cal E}_{1}$ produced in the reduction procedure retains competence property of its predecessor ${\bf E}$.
This result implies competence of matrix ${\bf D}_{n+1}$ in the factor $W({\bf D}_{n+1},{\bf s}-l{\bf c})$ in the convolution (68) derived in Section 5.1 and its further reduction to SPFs does not create any incompetent matrices. It means that any $l$-tuple partition of an incompetent $l$-row matrix eventually reduces to a superposition of scalar partition convolutions.
@@ 7 Conclusive remarks
The results presented in the manuscript show that Sylvester strongly underestimated the number of scalar partitions required for computation of a vector partition – their minimal number in case of $l\times m$ matrix significantly surpasses $m!/(m-l+1)!$ which in its turn is $(l-1)!$ times larger than predicted by Sylvester in [10].
We address the case of collinear columns of the generator matrix in Section 5 and show that in this case a vector partition boils down to a convolution of scalar partitions. We demonstrate that the reduction procedure does not produce new collinear columns.

### v2

@@ Abstract
A vector (compound) partition problem asks for a number of nonnegative integer solutions to a system of several linear Diophantine equations with integer nonnegative coefficients. J.J. Sylvester put forward an idea of reduction of vector partition to a sum of regular (scalar) partitions. For the simplest case of two equations with positive coefficients A. Cayley performed a reduction of the corresponding double partition to a sum of scalar partitions using an algorithm subject to a set of conditions on the coefficients. We suggested [9] a modification of the original Cayley algorithm for the cases when these conditions are not satisfied. In this manuscript we generalize the modified Cayley algorithm to arbitrary number of the Diophantine equations to accomplish the Sylvester program of the vector partition reduction to a combination of scalar partitions.
@@ 1 Introduction
The summation is made over all prime roots of unity $\rho_{j}=\exp(2\pi in/j)$ for $n$ relatively prime to $j$ (including unity) and smaller than $j$. It is possible to express any Sylvester wave $W_{j}(s,{\bf d})$ and the scalar partition $W(s,{\bf d})$ as a finite sum of the Bernoulli polynomials of higher order [6].
@@
Several approaches were suggested for VPF computation including the method of residues [1, 13] and geometric decomposition into so called chambers [10] – regions in $l$-dimensional space where the partition function is characterized by a specific expression. The “calculator” computer algebra system was developed to perform derivation of the explicit expressions of vector partitions over chambers [4]. Vector partitions appear to be a powerful tool for the computation of the Kronecker coefficients arising as the structure constants in the decomposition of a tensor product of irreducible representations of the symmetric group into irreducible representations [5]. A method of VPF computation was suggested being a direct generalization of the approach developed in [6] where vector Bernoulli and Eulerian polynomials of higher order were introduced to find explicit expression for $W({\bf s},{\bf D})$ [7]. A drawback of this approach is that it does not provide any mechanism to define the chamber boundaries. As SPF computation requires just known functions (the Bernoulli polynomials of higher order) [6] it is promising to obtain a reduction method expressing a vector partition through a combination of scalar ones.
The first step in this direction was made in the 19th century by J.J. Sylvester [11, 12] who suggested an iterative procedure of reduction of a vector partition into a sum of SPFs; in [11] he gave an example of such a reduction for a double partition. Based on this approach A. Cayley developed an algorithm of double partition reduction and established a set of conditions of this method applicability [2]. Cayley showed that each column ${\bf c}_{i}$ of the two-row matrix ${\bf D}$ gives rise to a scalar partition and elements of this column must be relatively prime (we call it P-column). Another limitation of the Cayley method is that the columns ${\bf c}_{i}$ should represent noncollinear vectors. A short historical excurse into the (seemingly forgotten) ideas of Sylvester and Cayley is given in Section 2 while the essence of the original Cayley algorithm is presented in Section 3.1.
@@
In Section 4 we extend the reduction algorithm to $l$-tuple (for arbitrary $l\geq 2$) vector partition defined by a $l\times m$ nonnegative generator matrix ${\bf D}$ made of $m$ columns ${\bf c}_{i}$. The first step of the algorithm discussed in Section 4.1 is the partial fraction expansion of the generating function (6) into $m$ terms corresponding to each column. The contribution of a column is computed in Section 4.2 and it is either a single partition for P-columns (Section 4.2.1) or a superposition of such partitions (Section 4.2.2) for NP-column. Three examples of triple partition reduction are given in Appendices.
The case of collinear columns of the generator matrix ${\bf D}$ is addressed in Section 5 where we show that in this case a vector partition reduces to a convolution of scalar partitions. Other details of computational implementation of the reduction step are presented in Sections 6.1 and 6.2 for P-column and NP-column respectively. We show in Section 6.3 that the reduction of a triple partition always produces NP-columns. This implies that the reduction of any VPF for a matrix with $l>2$ rows in most of the cases leads to application of the extended version of the Cayley algorithm. Additional details of the iterative reduction procedure are discussed in Section 6.4 while the concluding remarks are presented in Section 7.
@@ 2 Sylvester-Cayley method
Both manuscripts [11] and [2] were mentioned by L.E. Dickson in the then comprehensive volume [3] on the theory of numbers devoted to Diophantine analysis, but the literature search shows that both the reduction idea of Sylvester and the result of Cayley were forgotten. There are no publications discussing Sylvester research program of vector partition reduction to scalar ones. In my opinion the ideas and results of Sylvester and Cayley deserve not to be neglected, and this manuscript is an attempt to revive them and pave way to a general algorithm for vector partition expression through scalar partitions only.
@@ 3.4 Zero in last row of matrix ${\bf D}$
The matrix $\bm{D}^{\prime}_{1}$ used in derivation of (30) is assumed to be free of zeros in its last row. When this matrix has another column ${\bf c}_{j}$ with second zero element the columns ${\bf c}_{1}$ and ${\bf c}_{j}$ are parallel to each other and we refer the reader to Section 5 that considers the general case of collinear columns.
@@ 4.2.1 Compact form
It should be reminded here that the final goal of the Sylvester program is the VPF reduction to a set of SPFs and thus the elimination of a single (last) row of $l\times m$ matrix is just the first step of the procedure. We discuss the computation aspects of the Sylvester program further in Section 6.4 but now we have to underline that the reduction procedure described above is valid only when the matrix under reduction has last row void of nonpositive elements. It follows from (50) that some columns ${\bf c}^{1}_{j}$ might have negative elements $c^{1}_{jl}<0$ in the last position.
@@ 5 Vector partition with collinear columns
Call an integer generator matrix competent if its columns (viewed as vectors) are noncollinear and incompetent otherwise. Assume that vectors corresponding to the first $n$ columns of the incompetent generator matrix ${\bf D}$ are parallel to each other and rewrite the linear system ${\bf D}\cdot{\bf x}={\bf s}$ as follows
@@
where $u_{j}$ are positive integers and vector ${\bf C}$ represents the GCD vector of the collinear vectors ${\bf c}_{i},1\leq i\leq n<m$. The problem (66) is equivalent to
${\bf S}(k)={\bf s}-k{\bf C}=\sum_{i=n+1}^{m}{\bf c}_{i}x_{i},\quad k=\sum_{i=1}^{n}u_{i}x_{i},\quad 0\leq k\leq K=\min(s_{r}/C_{r}),\quad 1\leq r\leq l,$ (67)
@@
where $K$ is computed using only positive $C_{i}>0$. Introduce a vector ${\bf u}=\{u_{1},u_{2},\ldots,u_{n}\}$ and a competent matrix with $(m-n)$ columns $\bar{\bf D}=\{{\bf c}_{n+1},\ldots,{\bf c}_{m}\},$ for which the corresponding double partition $W({\bf S}(k),\bar{\bf D})$ admits a reduction to SPFs either as a sum of $W_{i}^{l}$ or a mixture of $W^{l}_{i}$ and $\bar{W}^{l}_{i}$. The number of solutions of the system (67) for given value of $k$ is equal to a number of solutions of the first equation given by the partition $W({\bf S}(k),\bar{\bf D})$ multiplied by a number of solutions of the second equation, i.e., by a scalar partition $W(k,{\bf u})$. Then the vector partition $W({\bf s},{\bf D})$ is equivalent to a convolution
$W({\bf s},{\bf D})=\sum_{k=0}^{K}W(k,{\bf u})H({\bf s}-k{\bf C})W({\bf s}-k{\bf C},\bar{\bf D}),\quad K=\min(s_{r}/C_{r}),\quad 1\leq r\leq l,$ (68)
of the scalar partition and the vector partition $W({\bf s}-k{\bf C},\bar{\bf D})$ built on the competent matrix $\bar{\bf D}$ (it admits the reduction described in Section 4). Note that the function $H({\bf s}-k{\bf C})$ in (68) is applied only to the elements of the vector ${\bf s}-k{\bf C}$ corresponding to the rows of $\bar{\bf D}$ void of negative generators.
This approach can be extended to the case when the matrix ${\bf D}$ has $r>1$ groups of collinear columns. Each $p$-th group is characterized by its GCD vector ${\bf C}_{p}$ and the $n_{p}$ integers $u_{p,i}$ making up the vector ${\bf u}_{p}=\{u_{p,1},u_{p,2},\ldots,u_{p,n_{p}}\}$. Introduce the vector ${\bf k}=\{k_{1},k_{2},\ldots,k_{r}\}$ and (67) turns into
${\bf S}({\bf k})={\bf s}-\sum_{p=1}^{r}k_{p}{\bf C}_{p}=\sum_{i=n+1}^{m}{\bf c}_{i}x_{i},\quad k_{p}=\sum_{i=1}^{n_{p}}u_{p,i}x_{i},\quad n=\sum_{p=1}^{r}n_{p}.$ (69)
The vector partition $W({\bf s},{\bf D})$ evaluates to the convolution
$W({\bf s},{\bf D})=\sum_{\bf k=0}^{\bf K}H({\bf S}({\bf k}))W({\bf S}({\bf k}),\bar{\bf D})\cdot\prod_{p=1}^{r}W(k_{p},{\bf u}_{p}),\ K_{p}=\min(s_{i}/C_{p,i}),\ \sum_{\bf k=0}^{\bf K}\equiv\sum_{k_{1}=0}^{K_{1}}\ldots\sum_{k_{r}=0}^{K_{r}}.$ (70)
Consider a special case $n=m-1$, i.e., when the competent matrix $\bar{\bf D}=\{{\bf c}\}$ has a single column ${\bf c}$. The vector partition $W({\bf S}({\bf k}),\{{\bf c}\})=1$ only when ${\bf S}({\bf k})=k_{0}{\bf c}$ for some nonnegative integer $k_{0}$, otherwise it vanishes. Thus we observe that $W({\bf S}({\bf k}),\bar{\bf D})=\delta_{{\bf S}({\bf k}),k_{0}{\bf c}}$ and we obtain using (70)
$W({\bf s},{\bf D})=\sum_{\bf k=0}^{\bf K}\sum_{k_{0}=0}^{K_{0}}\delta_{{\bf S}({\bf k}),k_{0}{\bf c}}H({\bf S}({\bf k}))\cdot\prod_{p=1}^{r}W(k_{p},{\bf u}_{p}),\quad K_{0}=\max(s_{i}/c_{i}),$ (71)
where again only positive $c_{i}$ are used for $K_{0}$ evaluation.
The extreme case arises for $n=m$, when each column of this matrix belongs to some group of the collinear columns. Use ${\bf k}^{\prime}=\{k_{1},k_{2},\ldots,k_{r-1}\}$ to write
${\bf S}({\bf k}^{\prime})={\bf s}-\sum_{p=1}^{r-1}k_{p}{\bf C}_{p}=k_{r}{\bf C}_{r}=\sum_{i=1}^{n_{r}}u_{r,i}x_{i}{\bf C}_{r}.$ (72)
As ${\bf S}({\bf k}^{\prime})$ and ${\bf C}_{r}$ are collinear we obtain a set of additional conditions on $k_{r}$ leading to
$W({\bf s},{\bf D})=\sum_{{\bf k}=0}^{{\bf K}}\delta_{{\bf S}({\bf k}^{\prime}),k_{r}{\bf C}_{r}}H({\bf S}({\bf k}^{\prime}))\cdot\prod_{p=1}^{r}W(k_{p},{\bf u}_{p}),\quad K_{r}=\max(s_{i}/C_{r,i}).$ (73)
The example of partitions with collinear columns is discussed in Appendix C.
@@ 7 Conclusive remarks
The results presented in the manuscript show that Sylvester strongly underestimated the number of scalar partitions required for computation of a vector partition – their minimal number in case of $l\times m$ matrix significantly surpasses $m!/(m-l+1)!$ which in its turn is $(l-1)!$ times larger than predicted by Sylvester in [12]. We address the case of collinear columns of the generator matrix in Section 5 and show that in this case a vector partition boils down to a convolution of scalar partitions.
@@ Appendix B Alternative triple partition reduction
Appendix C Partition with collinear columns
Consider reduction of a triple partition $W({\bf E})$ defined by the matrix
${\bf E}=\left(\begin{array}[]{cccccccccccc}s_{3}&0&0&1&1&1&1&2&1&2&2&3\\ s_{2}&0&1&0&0&1&1&1&1&1&2&2\\ s_{1}&1&0&0&1&0&0&0&1&1&1&1\end{array}\right).$ (C1)
It has two collinear columns $\{1,1,0\}^{T}$ but their contribution vanishes due to last zero element. As the last row of ${\bf E}$ has five zeros the first reduction step produces only six nonvanishing terms
$W({\bf E})=\sum_{k=1}^{6}A_{6}W({\bf E}_{k}),$ (C2)
where
$A_{1}=H(s_{2})H(s_{3}),$ $A_{2}=-H(s_{2}-2s_{1}-6)H(s_{3}-3s_{1}-9),$ (C3)
$A_{3}=-H(s_{2})H(s_{3}-s_{1}-1),$ $A_{4}=H(s_{2}-2s_{1}-6)H(s_{3}-2s_{1}-4),$
$A_{5}=H(s_{2}-s_{1}-2)H(s_{3}-s_{1}-1),$ $A_{6}=-H(s_{2}-s_{1}-2)H(s_{3}-2s_{1}-4),$
and the matrices ${\bf E}_{k}$ read
${\bf E}_{1}=\left(\begin{array}[]{ccccc|cc|cc|cc}s_{3}&1&1&1&2&2&2&1&1&3&0\\ s_{2}&1&1&1&2&1&1&0&0&2&1\end{array}\right),$ (C6)
${\bf E}_{2}=\left(\begin{array}[]{ccccc|cc|cc|cc}s_{3}-3s_{1}-9&1&1&1&2&2&2&1&1&3&0\\ s_{2}-2s_{1}-6&1&1&1&2&1&1&0&0&2&1\end{array}\right),$ (C9)
${\bf E}_{3}=\left(\begin{array}[]{ccccc|cc|cc|cc}s_{3}-s_{1}-1&1&1&1&2&0&0&1&1&1&2\\ s_{2}&1&1&1&2&1&1&0&0&2&1\end{array}\right),$ (C12)
${\bf E}_{4}=\left(\begin{array}[]{ccccc|cc|cc|cc}s_{3}-2s_{1}-4&1&1&1&2&0&0&1&1&1&2\\ s_{2}-2s_{1}-6&1&1&1&2&1&1&0&0&2&1\end{array}\right),$ (C15)
${\bf E}_{5}=\left(\begin{array}[]{ccccc|cc|cc|cc|}s_{3}-s_{1}-1&1&1&1&1&2&2&1&1&0&0\\ s_{2}-s_{1}-2&1&1&1&1&1&1&0&0&1&1\end{array}\right),$ (C18)
${\bf E}_{6}=\left(\begin{array}[]{ccccc|cc|cc|cc|}s_{3}-2s_{1}-4&1&1&1&1&2&2&1&1&0&0\\ s_{2}-s_{1}-2&1&1&1&1&1&1&0&0&1&1\end{array}\right),$ (C21)
where the groups of collinear columns are bounded from the right by a vertical line. Recalling that $W({\bf E}_{k})\equiv W({\bf s}_{k},{\bf D}_{k})$ we find three pairs of matrices with coinciding numerical parts ${\bf D}_{k}$ and different argument columns ${\bf s}_{k}$ – namely, ${\bf D}_{k}={\bf D}_{k+1}$ for $k=1,3,5$.
Use the recipies from Section 5 to rewrite $W({\bf s}_{1},{\bf D}_{1})$. We have ${\bf C}_{1}=\{1,1\}^{T}$ with ${\bf u}_{1}=\{1,1,1,2\}$; ${\bf C}_{2}=\{2,1\}^{T}$ with ${\bf u}_{2}=\{1,1\}$, and ${\bf C}_{3}=\{1,0\}^{T}$ with ${\bf u}_{3}=\{1,1\}$. Use (69) to write
${\bf S}_{1}({\bf k})={\bf s}_{1}-\sum_{p=1}^{3}k_{p}{\bf C}_{p}=\{s_{3}-k_{1}-2k_{2}-k_{3},s_{2}-k_{1}-k_{2}\}^{T},\quad\bar{\bf D}_{1}=\left(\begin{array}[]{cc}3&0\\ 2&1\end{array}\right),$ (C22)
and
$\!\!\!W({\bf s}_{1},{\bf D}_{1})\!=\!\!\!\sum_{k_{1}=0}^{K_{1}}\sum_{k_{2}=0}^{K_{2}}\sum_{k_{3}=0}^{K_{3}}H({\bf S}_{1}({\bf k}))W({\bf S}_{1}({\bf k}),\bar{\bf D}_{1})W(k_{1},\{1,1,1,2\})W(k_{2},\{1,1\})W(k_{3},\{1,1\}).$ (C23)
The upper limits $K_{i}$ in (C23) are determined by the nonnegativeness of the elements of the vector ${\bf S}_{1}({\bf k})$ in (C22). The relation (C23) admits further simplification as the scalar partition of any integer number $k$ into a set of $m$ units equals the binomial $\binom{k+m-1}{m-1}$, providing $W(k_{i},\{1,1\})=\binom{k_{i}+1}{1}=k_{i}+1$. The reduction of the double partition $W({\bf S}_{1}({\bf k}),\bar{\bf D}_{1})$ leads to
$W({\bf S}_{1}({\bf k}),\bar{\bf D}_{1})=W(s_{3}-k_{1}-2k_{2}-k_{3},\{3\})-W(2s_{3}-3s_{2}+k_{1}-k_{2}-k_{3}-3,\{3\}),$ (C24)
where both scalar partitions have the form $W(s,\{n\})$ that evaluates to $W(s,\{n\})=\delta_{s\bmod{n},0}$ for nonnegative $s$. Collecting these results we reduce (C23) to the following combination of the scalar partitions
$W({\bf s}_{1},{\bf D}_{1})$ $=$ $\sum_{k_{1}=0}^{K_{1}}\sum_{k_{2}=0}^{K_{2}}\sum_{k_{3}=0}^{K_{3}}H({\bf S}_{1}({\bf k}))(k_{2}+1)(k_{3}+1)W(k_{1},\{1,1,1,2\})$ (C25)
$\times$ $\left[W(s_{3}-k_{1}-2k_{2}-k_{3},\{3\})-W(2s_{3}-3s_{2}+k_{1}-k_{2}-2k_{3}-3,\{3\})\right].$
The term $W({\bf s}_{2},{\bf D}_{2})=W({\bf s}_{2},{\bf D}_{1})$ produces the relation similar to (C23) with the only change ${\bf S}_{1}({\bf k})\Rightarrow{\bf S}_{2}({\bf k})$ as follows
${\bf S}_{2}({\bf k})={\bf s}_{2}-\sum_{p=1}^{3}k_{p}{\bf C}_{p}=\{s_{3}-3s_{1}-9-k_{1}-2k_{2}-k_{3},s_{2}-2s_{1}-6-k_{1}-k_{2}\}^{T},$ (C26)
leading to
$W({\bf S}_{2}({\bf k}),\bar{\bf D}_{2})=W(s_{3}-3s_{1}-k_{1}-2k_{2}-k_{3},\{3\})-W(2s_{3}-3s_{2}+k_{1}-k_{2}-k_{3}-3,\{3\}).$ (C27)
Turn to evaluation of $W({\bf s}_{3},{\bf D}_{3})$ that is similar to $W({\bf s}_{1},{\bf D}_{1})$. We have ${\bf C}_{1}=\{1,1\}^{T}$ with ${\bf u}_{1}=\{1,1,1,2\}$; ${\bf C}_{2}=\{0,1\}^{T}$ with ${\bf u}_{2}=\{1,1\}$, and ${\bf C}_{3}=\{1,0\}^{T}$ with ${\bf u}_{3}=\{1,1\}$. Use (69) to write
${\bf S}_{3}({\bf k})={\bf s}_{3}-\sum_{p=1}^{3}k_{p}{\bf C}_{p}=\{s_{3}-s_{1}-1-k_{1}-k_{3},s_{2}-k_{1}-k_{2}\}^{T},$ (C28)
and
$W({\bf s}_{3},{\bf D}_{3})=\sum_{{\bf k}=0}^{\bf K}(k_{2}+1)(k_{3}+1)H({\bf S}_{3}({\bf k}))W({\bf S}_{3}({\bf k}),\bar{\bf D}_{3})W(k_{1},\{1,1,1,2\}),\ \bar{\bf D}_{3}=\left(\begin{array}[]{cc}1&2\\ 2&1\end{array}\right),$ (C29)
where the double partition $W({\bf S}_{3}({\bf k}),\bar{\bf D}_{3})$ reduces to
$W({\bf S}_{3}({\bf k}),\bar{\bf D}_{3})$ $=$ $W(2s_{3}-2s_{1}-s_{2}-2-k_{1}+k_{2}-2k_{3},\{3\})$ (C30)
-- $W(s_{3}-2s_{2}-s_{1}+k_{1}+2k_{2}-k_{3}-4,\{3\}).$
Similarly we obtain
$W({\bf S}_{4}({\bf k}),\bar{\bf D}_{4})$ $=$ $W(2s_{3}-2s_{1}-s_{2}-2-k_{1}+k_{2}-2k_{3},\{3\})$ (C31)
-- $W(s_{3}-2s_{2}+2s_{1}+k_{1}+2k_{2}-k_{3}+5,\{3\}),$
that replaces $W({\bf S}_{3}({\bf k}),\bar{\bf D}_{3})$ in (C29) to find explicit expression for $W({\bf s}_{4},{\bf D}_{4})$. Finally, consider $W({\bf s}_{5},{\bf D}_{5})$ with the matrix ${\bf D}_{5}$ void of noncollinear columns. This case is served by the relations (72,73) derived in Section 5. We choose the vectors ${\bf C}_{p}$ as follows ${\bf C}_{1}=\{2,1\}^{T}$ with ${\bf u}_{1}=\{1,1\}$; ${\bf C}_{2}=\{1,0\}^{T}$ with ${\bf u}_{2}=\{1,1\}$; ${\bf C}_{3}=\{0,1\}^{T}$ with ${\bf u}_{3}=\{1,1\}$, and ${\bf C}_{4}=\{1,1\}^{T}$ with ${\bf u}_{4}=\{1,1,1,1\}$. Use the vectors ${\bf s}_{5}=\{s_{3}-s_{1}-1,s_{2}-s_{1}-2\},\ {\bf s}_{6}=\{s_{3}-2s_{1}-4,s_{2}-s_{1}-2\}$ to find
${\bf S}_{5}({\bf k}^{\prime})={\bf s}_{5}-\sum_{p=1}^{3}k_{p}{\bf C}_{p}=\{s_{3}-s_{1}-1-2k_{1}-k_{2},s_{2}-s_{1}-2-k_{1}-k_{3}\}=\{k_{4},k_{4}\},$ (C32)
and the similar relation for ${\bf S}_{6}({\bf k}^{\prime})$. The expression (73) turns into
$W({\bf s}_{5},{\bf D}_{5})\!\!\!\!\!$ $=$ $\!\!\!\!\!\sum_{{\bf k}=0}^{{\bf K}}H({\bf S}_{5}({\bf k}^{\prime}))\binom{k_{4}+3}{3}\delta_{s_{3}-s_{1}-1-2k_{1}-k_{2},k_{4}}\delta_{s_{2}-s_{1}-2-k_{1}-k_{3},k_{4}}\prod_{j=1}^{3}(k_{j}+1),$ (C33)
$W({\bf s}_{6},{\bf D}_{6})\!\!\!\!\!$ $=$ $\!\!\!\!\!\sum_{{\bf k}=0}^{{\bf K}}H({\bf S}_{6}({\bf k}^{\prime}))\binom{k_{4}+3}{3}\delta_{s_{3}-2s_{1}-4-2k_{1}-k_{2},k_{4}}\delta_{s_{2}-s_{1}-2-k_{1}-k_{3},k_{4}}\prod_{j=1}^{3}(k_{j}+1).$ (C34)